# ShapeICP: Iterative Category-Level Object Pose and Shape Estimation from Depth

### v2

@@ III-D Coping with Local Minima
Figure 3: An occluded depth observation of the laptop (whose point cloud is fragmented in the right picture) causes ambiguity. The fitted model (whose vertices are visualized as the organized point set) has pose errors and collides with another object.
@@
Multi-hypothesis estimation. Initialization is critical in avoiding local minima. The translation can be initialized as the centroid of the observed point cloud $\bar{q}=\frac{1}{M}\sum_{m}{q_{m}}$. We find it better to begin with a small initial scale such as the average distance to the center $\frac{1}{M}\sum_{m}{q_{m}-\bar{q}}$. Without any neural network, the shape code is initialized to be the mean code for the category $\bar{\mathbf{c}}=\frac{1}{U}\sum_{u}{\tilde{\mathbf{c}}_{u}}$ where $U$ is the total number of database object models for the category. However, the rotation initial guess is not readily available. We thus track multiple rotation hypotheses in parallel and drop unpromising hypotheses quickly to save computation. Specifically, we start with the base $\mathrm{SO}(3)$ grid from [38], as also used by [26]. The grid has 2304 discrete rotations covering $\mathrm{SO}(3)$. During the course of optimization, we drop hypotheses according to the following score functions:
Objective function. The mean residual in (7) $\mathcal{S}_{r}=\frac{1}{M}\sum_{m}{r_{m}}=\frac{1}{M}\sum_{m}{\min_{n}{\|\hat{s}\hat{R}\hat{p}_{n}+\hat{t}-q_{m}\|_{2}^{2}}}$ plus the standard deviation of the residuals $\mathcal{S}_{\sigma}=\mathsf{std}(\{r_{m}\}_{m=1}^{M})$ is used. We find that adding the standard deviation improves the results, as we hope to have relatively uniform residuals over all the depth points.
@@
where $D$ is the masked observed depth image, $d\in\Omega(D)$ indexes pixels in the image space, $\mathsf{R}$ is the renderer, $\mathcal{M}(\hat{\mathbf{c}})$ is the mesh model built from $\hat{\mathbf{c}}$, and $K_{c}$ includes the known camera intrinsics and other camera parameters (such as image size). We use the renderer from [33] and assign the same background value to the rendered image and the masked observed depth image. The depth rendering helps to identify an over-sized model (which may still give low $\mathcal{S}_{r}$ if a portion of it fits the measurements well) [39]. To avoid computation overflow, we start this depth rendering after the number of hypotheses is low.
The total score to be minimized is $\mathcal{S}_{tot}=\mathcal{S}_{r}+\mathcal{S}_{\sigma}+\lambda_{\Psi}\mathcal{S}_{\Psi}+\lambda_{dr}\mathcal{S}_{dr}$ where $\lambda$’s are the weights. When picking top hypotheses, we greedily select the next hypothesis at least some angle ($\theta$) apart from the last chosen hypothesis to alleviate duplicated or very close hypotheses.
Shape classification. To better initialize the shape code $\mathbf{c}$, we build a shape classification network whose input is a color-coded normal vector image of the segmented object and the output is the index of the closest database model $u^{*}$. We choose ResNet-50 [40] for the network. The normal vector image is computed from the observed depth image and converted from $x,y,z\in[0,1]$ (assuming a unit vector) to RGB values. Only the image patch that contains the object (i.e. the 2D bounding box) is resized to a constant size and fed to the network. We find the ground-truth closest model by computing Chamfer distance (1) between the ground-truth shape and all the models in the database (ShapeNet [35]). The final initial guess is $\tilde{\mathbf{c}}_{u^{*}}$ for the shape code. Note that if a model in the database is not matched to any ground-truth shape, it does not appear as a class in the network output. Therefore, the number of classes is smaller than or equal to the number of ground-truth shapes.
@@ III-E Summary
             pick top $\text{pick top hypotheses}~{}\{\hat{R}_{g},\hat{t}_{g},\hat{s}_{g},\hat{\mathbf{c}}_{g}\}_{g=1}^{G_{new}}$;
@@
                   find the $\text{find the }Q\text{ closest points }\mathcal{N}_{Q}(q_{m})~{}\forall~{}q_{m}$;
[…]
                   /* (8) instead of (7) in $\mathcal{L}_{sc}$/
@@ IV-A Experiment Setup
Implementation. We implement our ShapeICP method with PyTorch [41] and PyTorch3D [33]. For the segmented depth image of an object, we perform statistical outlier removal on the back-projected point cloud as in [6, 42]. Moreover, we discard object detections that have fewer than a threshold number of points. All the hyperparameter values are summarized in the Appendix.
@@
Baselines. We benchmark our ShapeICP method against several interesting baselines. NOCS [21] is the seminal work in this research area. Neural Analysis-by-Synthesis [9] is an optimization-based method that is not learning-based and does not rely on image sequences or strong application-specific constraints (see Section II-A for more details). Metric Scale [11] is a rare case of adopting a mesh-based shape representation. ASM-Net [6] also uses ASM, but it is point-based. Unlike our method, its inference is mainly performed by neural network regression. Shape Prior [1] also estimates the shape by deforming a prior, but the method is learning-based. CenterSnap [22] is a one-shot end-to-end learning-based solution, which is the opposite of our work.
Evaluation protocol and metrics. We use the original public NOCS evaluation code for the pose evaluation, as was done in most prior work. A small bug in the NOCS code was later reported by [8]. We provide the updated results with the bug fixed in the Appendix. We use the Chamfer distance as the metric for the shape evaluation and follow the evaluation code from [1]. All the pose accuracy metrics in Tables I and III are in terms of the percentage of the results that fall under the specified thresholds. IoU is the 3D overlap between the estimated 3D bounding box and the ground-truth 3D bounding box of the object. Most neural methods directly predict three scales along the three axes ($x,y,z$) for more accurate 3D bounding box prediction. To build the 3D bounding box, our method multiplies the estimated scale by the aspect ratios along the three axes of the estimated shape, whose diagonal is normalized to unity (the final estimated scale absorbs the normalization factor).
@@ IV-B Results
TABLE I: Benchmarking results on NOCS REAL [21]. Baseline results are from sources [23, 6]. S.D. stands for shape-annotated data required for training and P.D. stands for pose-annotated data required for training. S.C. stands for the shape classification network. The results are roughly in the order of overall performance. The best result in each metric is in bold.
Method P.D. S.D. IoU25 IoU50 IoU75 5°5cm 5°10cm 10°5cm 10°10cm
NOCS [21] 84.8 78.0 30.1 10.0 9.8 25.2 25.8
Synthesis [9] 0.9 1.4 2.4 5.5
Metric Scale [11] 81.6 68.1 5.3 5.5 24.7 26.5
ASM-Net w/o ICP [6] 64.4 31.7 12.4 37.7
ShapeICP(Ours) w/o S.C. 82.6 58.3 37.0 32.4 33.2 42.2 43.4
ASM-Net w/ ICP [6] 68.3 37.4 25.6 43.7
ShapeICP(Ours) w/ S.C. 82.1 58.0 42.2 36.5 37.1 50.6 51.4
ShapePrior [1] 81.2 77.3 53.2 21.4 21.4 54.1 54.1
CenterSnap [22] 83.5 80.2 27.2 29.2 58.8 64.4
TABLE II: Shape estimation results on NOCS REAL [21] evaluated with Chamfer distance ($10^{-3}$). Baseline results are from sources [23, 6]. See the Appendix for per-class results.
@@
We emphasize the comparison between our method and ASM-Net [6] since we both adopt ASM. Despite not using any pose-annotated data, our method surpasses ASM-Net, which fully leverages all the annotated data, by a significant margin. Our no-learning variation (w/o S.C.) is even stronger than their ICP-assisted version in the 5°5cm metric and closely comparable in the IoU75 and 10°5cm metrics. Moreover, we attain a dominant win against the previous no-learning method, Neural Sythesis-by-Analysis [9], and the previous mesh-based method [11].
@@
In Fig. 4, we visualize some example results. The estimated objects in general fit the depth measurements very well, showing the effectiveness of ShapeICP in minimizing the objective function (7). The errors mostly stem from the inherent ambiguity in the problem. For example, a can is easily fit upside down. In addition, our method notably tends to make mistakes on mugs and cameras. The orientation of these objects mostly depends on local features such as the mug handle. Therefore, the global minimum for the orientation has a very small convergence basin, which is extremely difficult to find under the influence of noise and partial view. In general, our algorithm has a tendency to fit the bulk of the object well but ignore the local details.
The current focus of our algorithm is on accuracy, but a discussion of runtime is in the Appendix.
@@ IV-C Ablation
We ablate the modules for local minimum handling introduced in Section III-D in Table III. The symmetry check provides uniform improvements across the metrics. The EM treatment and the depth rendering check lead to more notable improvements in the higher accuracy metrics, such as 5°5cm and 5°2cm. The proposed multi-hypothesis estimation significantly outperforms single-hypothesis estimation that uses $\mathcal{S}_{r}+\mathcal{S}_{\sigma}$ to select the best hypothesis at the beginning of the iterations. More analysis is in the Appendix.
@@ V Conclusion
We proposed the ShapeICP algorithm for category-level object pose and shape estimation from a single depth image. Our method implements a novel mesh-based active shape model (ASM) to represent categorical object shapes. Remarkably, the method does not require pose-annotated data, unlike the vast majority of the methods in the field. However, the performance of ShapeICP surpasses or is on par with many learning-based methods that rely on pose-annotated data. Future directions include how to better disambiguate the problem, potentially by exploring RGB-image-based constraints, and how to better initialize the algorithm closer to the convergence basin of a global minimum.

### v3

@@ III-D Coping with Local Minima
Figure 3: The back-projected depths (blue points) are fragmented due to a partial view and occlusion, which causes ambiguity such that the models (red) fit the measurements well but have pose errors (in collision with each other).
@@
Multi-hypothesis estimation. Initialization is critical in avoiding local minima. The translation can be initialized as the centroid of the observed point cloud $\bar{q}=\frac{1}{M}\sum_{m}{q_{m}}$. We find it better to begin with a small initial scale such as the average distance to the center $\frac{1}{M}\sum_{m}{q_{m}-\bar{q}}$. Without any neural network, the shape code is initialized to be the mean code for the category $\bar{\mathbf{c}}=\frac{1}{U}\sum_{u}{\tilde{\mathbf{c}}_{u}}$ where $U$ is the total number of database object models for the category. However, the rotation initial guess is not readily available. We thus track multiple rotation hypotheses in parallel and drop unpromising hypotheses quickly to save computation. Specifically, we start with the base $\mathrm{SO}(3)$ grid from [38], as is also used by [26]. The grid has 2304 discrete rotations covering $\mathrm{SO}(3)$. During the course of optimization, we drop hypotheses according to the following score functions:
Objective function. The mean residual in (7) $\mathcal{S}_{r}=\frac{1}{M}\sum_{m}{r_{m}}=\frac{1}{M}\sum_{m}{\min_{n}{\|\hat{s}\hat{R}\hat{p}_{n}+\hat{t}-q_{m}\|_{2}^{2}}}$ plus the standard deviation of the residuals $\mathcal{S}_{\sigma}=\mathsf{std}(\{r_{m}\}_{m=1}^{M})$is used. We find that adding the standard deviation improves the results, as we hope to have relatively uniform residuals over all the depth measurement points.
@@
[…]
The total score to be minimized is $\mathcal{S}_{tot}=\mathcal{S}_{r}+\mathcal{S}_{\sigma}+\lambda_{\Psi}\mathcal{S}_{\Psi}+\lambda_{dr}\mathcal{S}_{dr}$ where the $\lambda$’s are the weights. When picking top hypotheses, we greedily select the next hypothesis at least some angle ($\theta$) apart from the last chosen hypothesis to alleviate duplicated or very close hypotheses.
Shape classification. To better initialize the shape code $\mathbf{c}$, we build a shape classification network whose input is a color-coded normal vector image of the segmented object and the output is the index of the closest database model $u^{*}$. We choose ResNet-50 [40] for the network. The normal vector image is computed from the observed depth image and converted from $x,y,z\in[0,1]$ (assuming a unit vector) to RGB values. Only the image patch that contains the object (i.e. the 2D bounding box) is resized to a constant size and fed to the network. We find the ground-truth closest model by computing the Chamfer distance (1) between the ground-truth shape and all the models in the database (ShapeNet [35]). The final initial guess is $\tilde{\mathbf{c}}_{u^{*}}$ for the shape code. Note that if a model in the database is not matched to any ground-truth shape, it does not appear as a class in the network output. Therefore, the number of classes is smaller than or equal to the number of ground-truth shapes.
@@ III-E Summary
     $\text{pick top hypotheses}~\{\hat{R}_{g},\hat{t}_{g},\hat{s}_{g},\hat{\mathbf{c}}_{g}\}_{g=1}^{G_{new}}$;
@@
        $\text{find the }Q\text{ closest points }\mathcal{N}_{Q}(q_{m})~\forall~q_{m}$;
        $\hat{\mathbf{c}}_{g}\leftarrow\text{minimize }\mathcal{L}_{sc}~\text{by one step of SGD}$;
        /* (8) instead of (7) in $\mathcal{L}_{sc}$ */
@@ IV-A Experiment Setup
Implementation. We implement our ShapeICP method with PyTorch [41] and PyTorch3D [33]. For the segmented depth image of an object, we perform statistical outlier removal on the back-projected point cloud as in [6, 42]. Moreover, we discard object detections that have fewer than a threshold number of points. All the hyperparameter values are given in the Appendix.
@@
Baselines. We benchmark our ShapeICP method against several interesting baselines. NOCS [21] is the seminal work in this research area. Neural Analysis-by-Synthesis [9] is an optimization-based method that is not learning-based for the purpose of pose inference and does not rely on image sequences or strong application-specific constraints (see Section II-A). Metric Scale [11] is a rare case of adopting a mesh-based shape representation. ASM-Net [6] also uses ASM, but it is point-based. Unlike our method, its inference is mainly performed by neural network regression. Shape Prior [1] also estimates the shape by deforming a prior, but the method is learning-based (i.e. regression-based). CenterSnap [22] is a one-shot end-to-end learning-based solution, which is the opposite of our work.
Evaluation protocol and metrics. We use the original public NOCS evaluation code for the pose evaluation, as is done in most prior work. A small bug in the NOCS code is later reported by [8]. We provide the updated results with the bug fixed in the Appendix. We use the Chamfer distance as the metric for the shape evaluation and follow the evaluation code from [1]. All the pose accuracy metrics in Tables I and III are in terms of the percentage of the results that fall under the specified thresholds. IoU is the 3D overlap between the estimated 3D bounding box and the ground-truth 3D bounding box of the object. Most neural methods directly predict three scales along the three axes ($x,y,z$) for more accurate 3D bounding box prediction. To build the 3D bounding box, our method multiplies the estimated scale by the aspect ratios along the three axes of the estimated shape, whose diagonal is normalized to one (the final estimated scale absorbs the normalization factor).
@@ IV-B Results
TABLE I: Benchmarking results on NOCS REAL [21]. Baseline results are from sources [23, 6]. P.D. stands for pose-annotated data required for training and S.D. stands for shape-annotated image data (but not including a database of categorical shape models). Inputs are the main inputs for inference. S.C. means shape classification. The results are roughly in the order of overall performance. The best result in each metric is in bold. 1 pose annotated images are required but synthetically generated 2 depth value for one pixel per object
Method P.D. S.D. Inputs IoU25 IoU50 IoU75 5°5cm 5°10cm 10°5cm 10°10cm
NOCS [21] RGB-D 84.8 78.0 30.1 10.0 9.8 25.2 25.8
Synthesis [9] 1 RGB 0.9 1.4 2.4 5.5
Metric Scale [11] RGB-OD2 81.6 68.1 5.3 5.5 24.7 26.5
ASM-Net w/o ICP [6] 1 Depth 64.4 31.7 12.4 37.7
ShapeICP(Ours) w/o S.C. Depth 82.6 58.3 37.0 32.4 33.2 42.2 43.4
ASM-Net w/ ICP [6] 1 Depth 68.3 37.4 25.6 43.7
ShapeICP(Ours) w/ S.C. Depth 82.1 58.0 42.2 36.5 37.1 50.6 51.4
ShapePrior [1] RGB-D 81.2 77.3 53.2 21.4 21.4 54.1 54.1
CenterSnap [22] RGB-D 83.5 80.2 27.2 29.2 58.8 64.4
TABLE II: Shape estimation results on NOCS REAL [21] evaluated with the Chamfer distance ($10^{-3}$). The baseline results are from sources [23, 6]. See the Appendix for the per-class results.
@@
We emphasize the comparison between our method and ASM-Net [6] since we both adopt an ASM. Despite not using any pose-annotated data, our method surpasses ASM-Net, which leverages annotated data, by a significant margin. Our no-learning variation (w/o S.C.) is even stronger than their ICP-assisted version in the 5°5cm metric and closely comparable in the IoU75 and 10°5cm metrics. Moreover, we attain a dominant win against the previous no-learning method, Neural Sythesis-by-Analysis [9], and the previous mesh-based method, Metric Scale [11].
@@
In Fig. 4, we visualize some example results. The estimated objects in general fit the depth measurements very well, showing the effectiveness of ShapeICP in minimizing the objective function (7). The errors mostly stem from the inherent ambiguity in the problem. For example, a can is easily fit upside down. In addition, our method notably tends to make mistakes on mugs and cameras. The orientation of these objects mostly depends on local features such as the mug handle. Therefore, the global minimum for the orientation has a very small convergence basin, which is extremely difficult to find under the influence of noise and partial views. In general, our algorithm has a tendency to fit the bulk of the object well but ignore the local details. More explanations are in the Appendix.
The current focus of our method is on feasibility and accuracy, while a discussion of runtime is in the Appendix.
@@ IV-C Ablation
We ablate the modules for local minimum handling introduced in Section III-D in Table III. The symmetry check provides uniform improvements across the metrics. The EM treatment and the depth rendering check lead to more notable improvements in the higher accuracy metrics, such as 5°5cm and 5°2cm. The proposed multi-hypothesis estimation significantly outperforms single-hypothesis estimation that uses $\mathcal{S}_{r}+\mathcal{S}_{\sigma}$ to select the best hypothesis at the beginning of the iteration. More analysis is in the Appendix.
@@ V Conclusion
We proposed the ShapeICP algorithm for category-level object pose and shape estimation from a single depth image. Our method implements a novel mesh-based active shape model (ASM) to represent categorical object shape. Remarkably, the method does not require pose-annotated data, unlike the vast majority of the methods in the field. However, the performance of ShapeICP surpasses or is on par with many learning-based methods that rely on pose-annotated data. Future directions include how to better disambiguate the problem, potentially by exploring RGB-image-based constraints, and how to better initialize the algorithm closer to the convergence basin of a global minimum.